OpenStudy (anonymous):

Find the slope of the tangent line to xtany = y-1 at y=pi/4

OpenStudy (anonymous):

Can anyone show the steps as well?

OpenStudy (lgbasallote):

i hated calculus but...slope of tangent means double derivative right?

OpenStudy (anonymous):

naw... just derivative

OpenStudy (anonymous):

are you differentiating with respect to x or y?

OpenStudy (lgbasallote):

but tangent is derivative right @dpaInc and slope is also derivative....right? or did i recall wrong :/

OpenStudy (anonymous):

it's just the first derivative....

OpenStudy (anonymous):

tangent is the line that kisses the curve at one point derivative is the slope of that line

OpenStudy (anonymous):

Im not sure @Hermeezey but the answer is 2/6-pi

OpenStudy (anonymous):

you'll have to take the derivative implicitly since it is a hassle to do it explicitly..

OpenStudy (anonymous):

xtany = y-1 derive the whole thing with respect to x, then solve for dy/dx tany + xsec^2y (dy/dx) = (dy/dx) xsec^2y(dy/dx) - (dy/dx) = -tany (dy/dx)(xsec^2y - 1) = -tany (dy/dx) = (-tany)/(xsec^2y - 1)

OpenStudy (anonymous):

Solve for y when x = pi/4. Plug x and y into dy/dx = (-tany)/(xsec^2y - 1) to get the slope of the tangent at that point.

OpenStudy (anonymous):

 solve for y' here...